# L4: Low-Latency and Load-Balanced LLM Serving via Length-Aware Scheduling

### v1

@@ 1 Introduction
We argue that the remedy must occur above the instance level. We observe that the massive scale of MILS, while typically viewed as a source of complexity, actually offers the solution to this problem. By scheduling requests among global instances, L4 turns the scale of MILS from a management burden into an optimization opportunity. Guided by this insight and the strong backend preference for homogeneous lengths, we propose to restructure MILS scheduling around the principle of: “Global homogenization for local efficiency.” Instead of having every instance cope with global heterogeneity, we reshape the cluster into a length-aware pipeline where each instance (or group of instances) processes requests of similar length. As requests grow during decoding, they move downstream through the pipeline, so every instance naturally forms length-homogeneous batches without modifying its internal scheduler.
This paper presents L4, a decentralized, length-aware inter-instance scheduling system for MILS clusters. L4 plans and maintains a multi-stage pipeline over the sequence-length space, routes incoming requests to the appropriate stage, and migrates them as they grow. To adapt to workload dynamics, L4 performs: (1) Workload-aware stage boundary adjustment: neighboring instances exchange lightweight length summaries and recompute boundaries to minimize predicted latency. (2) Decentralized bid-ask scheduling: overloaded instance offer requests while underloaded peers bid to acquire them, equalizing intra-stage load without a central coordinator.
@@
(2) We propose a shift from local-only to global-cooperative scheduling. By enforcing global length-homogeneity, L4 allows unmodified local schedulers to operate near hardware limits.
(3) We implement L4 on vLLM and demonstrate substantial latency and throughput improvements over SOTA serving systems.
@@ 2 Background and Motivation
Modern LLM serving clusters execute requests in a multi-instance setting, where each instance performs iterative token-by-token decoding using continuous batching [yu2022orca, kwon2023vllm, sglang]. While this abstraction worked well for short-context workloads, emerging long-context applications introduce a fundamental tension between how requests arrive and how attention kernels execute. This section summarizes the serving model, highlights the kernel-level characteristics that make sequence-length heterogeneity harmful, and explains why existing schedulers cannot address the resulting inefficiency. We conclude with the design insight that motivates L4.
@@ 2.5 Insight and Challenges
Challenge 2: The sequence length distribution can drift at runtime and become misaligned with the stage partition, leading to suboptimal performance. Stage partitioning is performed based on long-term workload statistics. But workloads and sequence-length distributions evolve over time. The misalignment between the partition and the distribution leads to performance degradation, but frequently (re)performing stage partitioning incurs significant overhead. Stage ranges must adapt to the distribution without disrupting ongoing decoding. L4 collects runtime length distribution and refines stage ranges to optimize the service quality (§4.3).
Challenge 3: Loads of instances in the same stage could evolve to be imbalanced, and the system needs a scalable way to rebalance them. Intra-stage load imbalance stems from inherent request diversity, compounded by stale state information and the irregularity of inter-stage request migrations. Naïvely applying a central scheduler to balance and migrate requests introduces latency and scalability issues, canceling out its performance gains. L4 devises a decentralized bid-ask protocol: instances negotiate request migrations via sender-receiver bids, enabling efficient, scalable workload rebalancing (§4.4).
3 L4 Overview
@@ 3.1 Architecture
When serving an LLM, L4 organizes inference instances into multiple length-specialized groups, each dedicated to sequences within a specific length range. During autoregressive decoding, sequence lengths increase monotonically. L4 leverages this property to form a logical pipeline, with each group as a pipeline stage; as decoding progresses, sequences naturally flow from shorter to longer stages. As shown in Figure˜3, this pipeline comprises $N$ stages, each with a variable number of instances. Within each instance, L4 retains standard components: a scheduler for sequence batching and a GPU executor for batch processing on the attention backend. To support adaptive, decentralized rescheduling, L4 introduces two lightweight runtime components—LoadTracker and Coordinator—in each instance.
The LoadTracker monitors the instance’s token-level workload (e.g., the number of input and output tokens per request) and exchanges these statistics with other instances—both peers in the same stage and instances in adjacent stages. This real-time monitoring allows L4 to capture transient workload fluctuations and trigger responsive adjustments. The Coordinator manages dynamic sequence migration by transferring metadata and KV cache across instances, either within a stage or to the next stage. By operating in a decentralized manner, the Coordinator eliminates the bottleneck of a central scheduler, thereby ensuring cluster scalability and minimizing the risk of scheduling decisions based on stale state information.
@@ 3.2 Workflow
System bootup. L4 constructs the instance pipeline at startup. It applies a length-aware stage partition algorithm (§4.2) to partition instances into length-specialized stages. Each instance initializes its runtime components and establishes connections to peer instances within the same stage and to instances in adjacent stages. These connections enable both intra-stage and cross-stage coordination.
Runtime request processing. L4 performs progressive decoding with stage transitions as described in Section˜2.2. As shown in Figure˜3, when a request arrives, it is routed to the earliest stage whose serving range covers its initial length. The assigned instance adds this request to its batch when memory is available, then performs prefill and decoding locally using its scheduler and GPU executor. As the sequence grows, if its length exceeds the instance’s range, it is migrated to the next stage for further decoding. This process repeats until the sequence completes.
Figure 3: Architecture and workflow of L4. Engine instances are grouped by length into stages forming a logical pipeline; sequences may exit early without traversing all stages.
@@ 4.2 Length-Aware Stage Partition
Pipeline planning runs at initialization and periodically thereafter. L4 collects request statistics over the period (input and output lengths) and generates pipeline settings for the next period.
@@ 4.3 Adaptive Range Refinement
Each instance periodically refines its range boundary based on LoadTracker. At each refinement interval, the instance collects the sequence lengths from both itself and its successors. L4 first computes the average workload of the next stage: it merges the successor sequence lengths as a union set, and divides the set evenly by the number of successors (using the same set division method in Section˜4.2).
L4 merges the instance’s local sequence lengths and the average successor lengths into a single set. It sorts the requests in the set as a list $R$ with requests indexed from 0 to N1N-1. L4 enumerates all possible splitting points in $R$, computes the overall quality of both parts by Equation˜1, and finds the optimal one with the best overall quality:
@@ 4.4 Decentralized Load (Re)Balancing
Bid-ask protocol. For scalability, L4 employs a decentralized bid-ask scheduling protocol: each sender distributes its requests to receivers independently, and all senders’ requests are balanced across all receivers.
@@ 5 Implementation
L4 is implemented with  3,000 lines of Python for the control plane and 700 lines of C++ for the high-performance communication backend, built atop vLLM (v0.9.1) without modifying its core attention kernels. We integrate L4 modularly into vLLM via Scheduler-Aware Hooks: the LoadTracker monitors token-level workloads, while the Coordinator manages metadata and orchestrates KV cache migrations.
Efficient migration infrastructure. To minimize the overhead of frequent, fine-grained KV transfers, we implement a custom C++ communication backend that bypasses Python’s GIL and avoids the latency of NCCL’s collective operations, which are ill-suited for numerous small messages. Instead, the backend leverages low-level APIs—specifically cudaMemcpyPeerAsync for intra-node and RDMA verbs for inter-node transfers—to achieve zero-copy GPU-to-GPU transmission. Furthermore, L4 optimizes physical topology by placing instances of adjacent pipeline stages on the same node whenever possible, thereby significantly reducing expensive inter-node communication.
Migration flow control and optimization. L4 optimizes runtime execution by eliminating intermediate buffers; KV caches are transferred directly into idle slots on the target instance, and migration is skipped if no idle cache is available. To prevent bandwidth contention, L4 enforces a strict concurrency limit (capped at three parallel transfers in our implementation); requests exceeding this threshold continue running on the source to avoid performance regression. Finally, we employ asynchronous multi-round live migration (adapting Llumnix [sun2024llumnix]) combined with bidirectional transfer support and a starvation-aware backpressure mechanism. This design effectively overlaps communication with computation, preventing head-of-line blocking while maximizing GPU and NIC utilization.
@@ 6 Evaluation
(1) How does L4 affect request latency across diverse configurations? (Section˜6.2)
(2) How does L4 influence system throughput across diverse configurations? (Section˜6.3)
(3) How does L4 impact SLO attainment? (Section˜6.4)
(4) What is the effect of each L4 mechanism? (Section˜6.5)
@@ 6.1 Experimental Setup
Testbed. We evaluate L4 on two testbeds, each comprising two nodes with 16 NVIDIA GPUs. The first testbed uses NVLink-connected H20 GPUs, each with 141 GB of memory, while the second employs PCIe-connected L40 GPUs, each with 48 GB. Every node in both setups is equipped with 2 TB of host memory, an Intel® Xeon® Platinum 8468 CPU (192 logical cores), and a 400 Gbps ConnectX-7 NIC.
Baselines. We compare L4 against three popular open-source inference frameworks: vLLM 0.9.1 [kwon2023vllm], SGLang 0.4.9 [sglang], and Llumnix [sun2024llumnix]. vLLM and SGLang operate as standalone inference service instances without inter-instance scheduling; we therefore deploy a simple round-robin load balancer across their instances. In contrast, Llumnix extends vLLM with an integrated inter-instance scheduler, which we run using its default configuration. Unless stated otherwise, all backends use an FCFS policy. We cap the batch size at 1024—the vLLM official default—with actual batch sizes further constrained by memory and token budgets at runtime. Following the official default settings, we use FlashAttention [dao2022flashattention] backend for vLLM, Llumnix, and L4, and use FlashInfer [ye2025flashinfer] backend for SGLang [sglang].
Models. We evaluate eight LLMs in FP16 precision across a range of commonly used sizes, grouped into four categories: (1) Tiny — Llama-3.2-3B [llama-3.2] and Phi-3-3B [team2024phi]; (2) Small — Llama-3.1-8B [llama-3.1] and GLM-4-9B [glm2024chatglm]; (3) Moderate — Phi-3-14B [team2024phi] and Qwen-2.5-14B [qwen2025qwen25technicalreport]; and (4) Large — QwQ-32B [qwq] and Qwen-2.5-32B [qwen2025qwen25technicalreport]. Each instance runs on a single GPU, except for Llama-3.1-70B [llama-3.1], which we evaluate in 2- and 4-way tensor-parallel configurations. [^2] We make homogeneous instance setups and leave the discussion of heterogeneous configuration in Section 8. We exclude larger models because their high per-instance GPU requirements would leave too few instances on our testbeds, misrepresenting typical multi-instance deployments. L4 constructs pipelines of 4 to 6 stages for the selected model, with each stage comprising 1 to 4 instances. All selected models support context lengths of at least 128K tokens.
@@
Experiment parameters. We vary request arrival rates to cover both light and heavy loads. Light load verifies that L4 does not introduce a negative impact, while heavy load, our target scenario, evaluates its performance gains.
@@ 6.2 Reduced End-to-End Latency
Latency reduction across models. We compare L4 with other systems across eight LLMs on the H20 testbed. The results are shown in Figures 6 and 7.
Figure 8: TPOT of a single instance across varying request arrival rates. L4’s single-instance performance matches vLLM’s but falls behind Llumnix’s. By comparing this to other results in §6, we find that L4’s multi-instance scheduling delivers higher gains than Llumnix’s.
TTFT. As shown in Figure˜6, L4 consistently achieves a lower mean TTFT than vLLM, SGLang, and Llumnix. Under heavy load, it reduces mean TTFT by 67%–78% compared to vLLM, by 70%–84% compared to SGLang, and by 36%–66% compared to Llumnix, the state-of-the-art in multi-instance scheduling system. Tail TTFT exhibits similar trends, with average reductions of 56%, 62%, and 19%, respectively. Under light load, model access overhead (unrelated to requests) dominates latency, so the system reverts to standard multi-instance scheduling with performance comparable to Llumnix.
TPOT. As shown in Figure˜7, L4 achieves a lower mean TPOT than vLLM, SGLang, and Llumnix in most settings. Under heavy load, it reduces mean TPOT by 30%–64% vs. vLLM, 25%–77% vs. SGLang, and 3.4%–56% vs. Llumnix. L4 also reduces tail TPOT by 25% vs. vLLM, 61% vs. SGLang, and 20% vs. Llumnix under heavy load.
Under light load, Llumnix performs better because it runs on a newer engine with lower latency. Since L4 focuses on scheduling requests across engine instances rather than optimizing a single engine instance, its benefits here are actually underestimated—on that same newer engine, L4 would achieve even lower TPOT. We confirmed this by experiments below. The results show that L4’s single-instance performance is on par with vLLM’s but falls short of Llumnix’s.
@@
Comparison of single-instance performance. L4 focuses on improving cluster‐level performance by using length‐aware request scheduling across instances. When processing requests from the same batch, L4 delivers single‐instance performance on par with vLLM, demonstrating full compatibility with both existing and future intra‐instance optimizations. In a side-by-side comparison of single-instance performance among L4, vLLM, and Llumnix (Figure˜8), L4 matches vLLM but is still outperformed by Llumnix’s latest engine by 22%–81% since Llumnix’s engine has a lower runtime scheduling and executing overhead than vLLM. This gap indicates that the cluster‐level gains reported in the evaluation actually underestimate L4’s potential—when paired with the newest engine, L4 can achieve even higher overall performance.
Latency reduction across testbeds. To demonstrate that L4’s performance gains are platform-independent, we evaluate it on the L40 testbed, which has 48 GB of memory and thus only supports small models. We measure normalized latency, defined as the end-to-end delay per output token, including both TTFT and TPOT. As shown in Figure˜9(a), L4 outperforms all baselines and reduces latency by 45%–67%.
Latency reduction under tensor parallelism. To evaluate L4 on large models spanning multiple GPUs, we run Llama-3.1-70B with tensor parallelism (TP) of 2 and 4 on the H20 testbed. At TP=2, model parameters occupy nearly half of each GPU’s memory, causing higher parameter access latency and reducing L4’s effectiveness compared to TP=4. As shown in Figure˜9(b), L4 reduces latency by 27%–65% with 2 GPUs and 49%–64% with 4 GPUs, outperforming all baselines.

### v2

# CascadeInfer: Low-Latency and Load-Balanced LLM Serving via Length-Aware Scheduling

@@ 1 Introduction
We argue that the remedy must occur above the instance level. We observe that the massive scale of MILS, while typically viewed as a source of complexity, actually offers the solution to this problem. By scheduling requests among global instances, CascadeInfer turns the scale of MILS from a management burden into an optimization opportunity. Guided by this insight and the strong backend preference for homogeneous lengths, we propose to restructure MILS scheduling around the principle of: “Global homogenization for local efficiency.” Instead of having every instance cope with global heterogeneity, we reshape the cluster into a length-aware pipeline where each instance (or group of instances) processes requests of similar length. As requests grow during decoding, they move downstream through the pipeline, so every instance naturally forms length-homogeneous batches without modifying its internal scheduler.
This paper presents CascadeInfer, a decentralized, length-aware inter-instance scheduling system for MILS clusters. CascadeInfer plans and maintains a multi-stage pipeline over the sequence-length space, routes incoming requests to the appropriate stage, and migrates them as they grow. To adapt to workload dynamics, CascadeInfer performs: (1) Workload-aware stage boundary adjustment: neighboring instances exchange lightweight length summaries and recompute boundaries to minimize predicted latency. (2) Decentralized bid-ask scheduling: overloaded instance offer requests while underloaded peers bid to acquire them, equalizing intra-stage load without a central coordinator.
@@
(2) We propose a shift from local-only to global-cooperative scheduling. By enforcing global length-homogeneity, CascadeInfer allows unmodified local schedulers to operate near hardware limits.
(3) We implement CascadeInfer on vLLM and demonstrate substantial latency and throughput improvements over SOTA serving systems.
@@ 2 Background and Motivation
Modern LLM serving clusters execute requests in a multi-instance setting, where each instance performs iterative token-by-token decoding using continuous batching [yu2022orca, kwon2023vllm, sglang]. While this abstraction worked well for short-context workloads, emerging long-context applications introduce a fundamental tension between how requests arrive and how attention kernels execute. This section summarizes the serving model, highlights the kernel-level characteristics that make sequence-length heterogeneity harmful, and explains why existing schedulers cannot address the resulting inefficiency. We conclude with the design insight that motivates CascadeInfer.
@@ 2.5 Insight and Challenges
Challenge 2: The sequence length distribution can drift at runtime and become misaligned with the stage partition, leading to suboptimal performance. Stage partitioning is performed based on long-term workload statistics. But workloads and sequence-length distributions evolve over time. The misalignment between the partition and the distribution leads to performance degradation, but frequently (re)performing stage partitioning incurs significant overhead. Stage ranges must adapt to the distribution without disrupting ongoing decoding. CascadeInfer collects runtime length distribution and refines stage ranges to optimize the service quality (§4.3).
Challenge 3: Loads of instances in the same stage could evolve to be imbalanced, and the system needs a scalable way to rebalance them. Intra-stage load imbalance stems from inherent request diversity, compounded by stale state information and the irregularity of inter-stage request migrations. Naïvely applying a central scheduler to balance and migrate requests introduces latency and scalability issues, canceling out its performance gains. CascadeInfer devises a decentralized bid-ask protocol: instances negotiate request migrations via sender-receiver bids, enabling efficient, scalable workload rebalancing (§4.4).
3 CascadeInfer Overview
@@ 3.1 Architecture
When serving an LLM, CascadeInfer organizes inference instances into multiple length-specialized groups, each dedicated to sequences within a specific length range. During autoregressive decoding, sequence lengths increase monotonically. CascadeInfer leverages this property to form a logical pipeline, with each group as a pipeline stage; as decoding progresses, sequences naturally flow from shorter to longer stages. As shown in Figure˜3, this pipeline comprises $N$ stages, each with a variable number of instances. Within each instance, CascadeInfer retains standard components: a scheduler for sequence batching and a GPU executor for batch processing on the attention backend. To support adaptive, decentralized rescheduling, CascadeInfer introduces two lightweight runtime components—LoadTracker and Coordinator—in each instance.
The LoadTracker monitors the instance’s token-level workload (e.g., the number of input and output tokens per request) and exchanges these statistics with other instances—both peers in the same stage and instances in adjacent stages. This real-time monitoring allows CascadeInfer to capture transient workload fluctuations and trigger responsive adjustments. The Coordinator manages dynamic sequence migration by transferring metadata and KV cache across instances, either within a stage or to the next stage. By operating in a decentralized manner, the Coordinator eliminates the bottleneck of a central scheduler, thereby ensuring cluster scalability and minimizing the risk of scheduling decisions based on stale state information.
@@ 3.2 Workflow
System bootup. CascadeInfer constructs the instance pipeline at startup. It applies a length-aware stage partition algorithm (§4.2) to partition instances into length-specialized stages. Each instance initializes its runtime components and establishes connections to peer instances within the same stage and to instances in adjacent stages. These connections enable both intra-stage and cross-stage coordination.
Runtime request processing. CascadeInfer performs progressive decoding with stage transitions as described in Section˜2.2. As shown in Figure˜3, when a request arrives, it is routed to the earliest stage whose serving range covers its initial length. The assigned instance adds this request to its batch when memory is available, then performs prefill and decoding locally using its scheduler and GPU executor. As the sequence grows, if its length exceeds the instance’s range, it is migrated to the next stage for further decoding. This process repeats until the sequence completes.
Figure 3: Architecture and workflow of CascadeInfer. Engine instances are grouped by length into stages forming a logical pipeline; sequences may exit early without traversing all stages.
@@ 4.2 Length-Aware Stage Partition
Pipeline planning runs at initialization and periodically thereafter. CascadeInfer collects request statistics over the period (input and output lengths) and generates pipeline settings for the next period.
@@ 4.3 Adaptive Range Refinement
Each instance periodically refines its range boundary based on LoadTracker. At each refinement interval, the instance collects the sequence lengths from both itself and its successors. CascadeInfer first computes the average workload of the next stage: it merges the successor sequence lengths as a union set, and divides the set evenly by the number of successors (using the same set division method in Section˜4.2).
CascadeInfer merges the instance’s local sequence lengths and the average successor lengths into a single set. It sorts the requests in the set as a list $R$ with requests indexed from 0 to N1N-1. CascadeInfer enumerates all possible splitting points in $R$, computes the overall quality of both parts by Equation˜1, and finds the optimal one with the best overall quality:
@@ 4.4 Decentralized Load (Re)Balancing
Bid-ask protocol. For scalability, CascadeInfer employs a decentralized bid-ask scheduling protocol: each sender distributes its requests to receivers independently, and all senders’ requests are balanced across all receivers.
@@ 5 Implementation
CascadeInfer is implemented with  3,000 lines of Python for the control plane and 700 lines of C++ for the high-performance communication backend, built atop vLLM (v0.9.1) without modifying its core attention kernels. We integrate CascadeInfer modularly into vLLM via Scheduler-Aware Hooks: the LoadTracker monitors token-level workloads, while the Coordinator manages metadata and orchestrates KV cache migrations.
Efficient migration infrastructure. To minimize the overhead of frequent, fine-grained KV transfers, we implement a custom C++ communication backend that bypasses Python’s GIL and avoids the latency of NCCL’s collective operations, which are ill-suited for numerous small messages. Instead, the backend leverages low-level APIs—specifically cudaMemcpyPeerAsync for intra-node and RDMA verbs for inter-node transfers—to achieve zero-copy GPU-to-GPU transmission. Furthermore, CascadeInfer optimizes physical topology by placing instances of adjacent pipeline stages on the same node whenever possible, thereby significantly reducing expensive inter-node communication.
Migration flow control and optimization. CascadeInfer optimizes runtime execution by eliminating intermediate buffers; KV caches are transferred directly into idle slots on the target instance, and migration is skipped if no idle cache is available. To prevent bandwidth contention, CascadeInfer enforces a strict concurrency limit (capped at three parallel transfers in our implementation); requests exceeding this threshold continue running on the source to avoid performance regression. Finally, we employ asynchronous multi-round live migration (adapting Llumnix [sun2024llumnix]) combined with bidirectional transfer support and a starvation-aware backpressure mechanism. This design effectively overlaps communication with computation, preventing head-of-line blocking while maximizing GPU and NIC utilization.
@@ 6 Evaluation
(1) How does CascadeInfer affect request latency across diverse configurations? (Section˜6.2)
(2) How does CascadeInfer influence system throughput across diverse configurations? (Section˜6.3)
(3) How does CascadeInfer impact SLO attainment? (Section˜6.4)
(4) What is the effect of each CascadeInfer mechanism? (Section˜6.5)
@@ 6.1 Experimental Setup
Testbed. We evaluate CascadeInfer on two testbeds, each comprising two nodes with 16 NVIDIA GPUs. The first testbed uses NVLink-connected H20 GPUs, each with 141 GB of memory, while the second employs PCIe-connected L40 GPUs, each with 48 GB. Every node in both setups is equipped with 2 TB of host memory, an Intel® Xeon® Platinum 8468 CPU (192 logical cores), and a 400 Gbps ConnectX-7 NIC.
Baselines. We compare CascadeInfer against three popular open-source inference frameworks: vLLM 0.9.1 [kwon2023vllm], SGLang 0.4.9 [sglang], and Llumnix [sun2024llumnix]. vLLM and SGLang operate as standalone inference service instances without inter-instance scheduling; we therefore deploy a simple round-robin load balancer across their instances. In contrast, Llumnix extends vLLM with an integrated inter-instance scheduler, which we run using its default configuration. Unless stated otherwise, all backends use an FCFS policy. We cap the batch size at 1024—the vLLM official default—with actual batch sizes further constrained by memory and token budgets at runtime. Following the official default settings, we use FlashAttention [dao2022flashattention] backend for vLLM, Llumnix, and CascadeInfer, and use FlashInfer [ye2025flashinfer] backend for SGLang [sglang].
Models. We evaluate eight LLMs in FP16 precision across a range of commonly used sizes, grouped into four categories: (1) Tiny — Llama-3.2-3B [llama-3.2] and Phi-3-3B [team2024phi]; (2) Small — Llama-3.1-8B [llama-3.1] and GLM-4-9B [glm2024chatglm]; (3) Moderate — Phi-3-14B [team2024phi] and Qwen-2.5-14B [qwen2025qwen25technicalreport]; and (4) Large — QwQ-32B [qwq] and Qwen-2.5-32B [qwen2025qwen25technicalreport]. Each instance runs on a single GPU, except for Llama-3.1-70B [llama-3.1], which we evaluate in 2- and 4-way tensor-parallel configurations. [^2] We make homogeneous instance setups and leave the discussion of heterogeneous configuration in Section 8. We exclude larger models because their high per-instance GPU requirements would leave too few instances on our testbeds, misrepresenting typical multi-instance deployments. CascadeInfer constructs pipelines of 4 to 6 stages for the selected model, with each stage comprising 1 to 4 instances. All selected models support context lengths of at least 128K tokens.
@@
Experiment parameters. We vary request arrival rates to cover both light and heavy loads. Light load verifies that CascadeInfer does not introduce a negative impact, while heavy load, our target scenario, evaluates its performance gains.
@@ 6.2 Reduced End-to-End Latency
Latency reduction across models. We compare CascadeInfer with other systems across eight LLMs on the H20 testbed. The results are shown in Figures 6 and 7.
Figure 8: TPOT of a single instance across varying request arrival rates. CascadeInfer’s single-instance performance matches vLLM’s but falls behind Llumnix’s. By comparing this to other results in §6, we find that CascadeInfer’s multi-instance scheduling delivers higher gains than Llumnix’s.
TTFT. As shown in Figure˜6, CascadeInfer consistently achieves a lower mean TTFT than vLLM, SGLang, and Llumnix. Under heavy load, it reduces mean TTFT by 67%–78% compared to vLLM, by 70%–84% compared to SGLang, and by 36%–66% compared to Llumnix, the state-of-the-art in multi-instance scheduling system. Tail TTFT exhibits similar trends, with average reductions of 56%, 62%, and 19%, respectively. Under light load, model access overhead (unrelated to requests) dominates latency, so the system reverts to standard multi-instance scheduling with performance comparable to Llumnix.
TPOT. As shown in Figure˜7, CascadeInfer achieves a lower mean TPOT than vLLM, SGLang, and Llumnix in most settings. Under heavy load, it reduces mean TPOT by 30%–64% vs. vLLM, 25%–77% vs. SGLang, and 3.4%–56% vs. Llumnix. CascadeInfer also reduces tail TPOT by 25% vs. vLLM, 61% vs. SGLang, and 20% vs. Llumnix under heavy load.
Under light load, Llumnix performs better because it runs on a newer engine with lower latency. Since CascadeInfer focuses on scheduling requests across engine instances rather than optimizing a single engine instance, its benefits here are actually underestimated—on that same newer engine, CascadeInfer would achieve even lower TPOT. We confirmed this by experiments below. The results show that CascadeInfer’s single-instance performance is on par with vLLM’s but falls short of Llumnix’s.
@@
Comparison of single-instance performance. CascadeInfer focuses on improving cluster‐level performance by using length‐aware request scheduling across instances. When processing requests from the same batch, CascadeInfer delivers single‐instance performance on par with vLLM, demonstrating full compatibility with both existing and future intra‐instance optimizations. In a side-by-side comparison of single-instance performance among CascadeInfer, vLLM, and Llumnix (Figure˜8), CascadeInfer matches vLLM but is still outperformed by Llumnix’s latest engine by 22%–81% since Llumnix’s engine has a lower runtime scheduling and executing overhead than vLLM. This gap indicates that the cluster‐level gains reported in the evaluation actually underestimate CascadeInfer’s potential—when paired with the newest engine, CascadeInfer can achieve even higher overall performance.
Latency reduction across testbeds. To demonstrate that CascadeInfer’s performance gains are platform-independent, we evaluate it on the L40 testbed, which has 48 GB of memory and thus only supports small models. We measure normalized latency, defined as the end-to-end delay per output token, including both TTFT and TPOT. As shown in Figure˜9(a), CascadeInfer outperforms all baselines and reduces latency by 45%–67%.
Latency reduction under tensor parallelism. To evaluate CascadeInfer on large models spanning multiple GPUs, we run Llama-3.1-70B with tensor parallelism (TP) of 2 and 4 on the H20 testbed. At TP=2, model parameters occupy nearly half of each GPU’s memory, causing higher parameter access latency and reducing CascadeInfer’s effectiveness compared to TP=4. As shown in Figure˜9(b), CascadeInfer reduces latency by 27%–65% with 2 GPUs and 49%–64% with 4 GPUs, outperforming all baselines.